Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution.
Problem:
Chuck and Dana agree to meet in Chicago for the weekend. Chuck travels 114 miles in the same time Dana travels 96 miles. If Chuck's rate of travel is 6mph more that Dana's, then at what rate does Chuck travels?

Answer by mananth(16946)

Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution.
Problem:
Chuck and Dana agree to meet in Chicago for the weekend. Chuck travels 114 miles in the same time Dana travels 96 miles. If Chuck's rate of travel is 6mph more that Dana's, then at what rate does Chuck travels?
Let dana's speed be x mph
Chuck's rate will b x+6 mph
time taken by chuck = 114 / x+6
time taken by dana = 96/x
The times are same
114/ x+6 = 96/x
114/96 = x+6 / x
114x /96= x+6
114x = 96x+576
18x= 576
x=32
Chuck travels at 38 mph.